# Self-Play Fine-Tuning Converts Weak Language Models to Strong Language Models

### 02 Jan 2024 v1

@@ Abstract
Harnessing the power of human-annotated data through Supervised Fine-Tuning (SFT) is pivotal for advancing Large Language Models (LLMs). In this paper, we delve into the prospect of growing a strong LLM out of a weak one without the need for acquiring additional human-annotated data. We propose a new fine-tuning method called Self-Play fIne-tuNing (SPIN), which starts from a supervised fine-tuned model. At the heart of SPIN lies a self-play mechanism, where the LLM refines its capability by playing against instances of itself. More specifically, the LLM generates its own training data from its previous iterations, refining its policy by discerning these self-generated responses from those obtained from human-annotated data. Our method progressively elevates the LLM from a nascent model to a formidable one, unlocking the full potential of human-annotated demonstration data for SFT. Theoretically, we prove that the global optimum to the training objective function of our method is achieved only when the LLM policy aligns with the target data distribution. Empirically, we evaluate our method on several benchmark datasets including the HuggingFace Open LLM Leaderboard, MT-Bench, and datasets from Big-Bench. Our results show that SPIN can significantly improve the LLM’s performance across a variety of benchmarks and even outperform models trained through direct preference optimization (DPO) supplemented with extra GPT-4 preference data. This sheds light on the promise of self-play, enabling the achievement of human-level performance in LLMs without the need for expert opponents.
@@ 1 Introduction
Notation. We use lowercase letters and lowercase boldface letters to denote scalars and vectors, respectively. We use $[N]$ to denote the index set $\{1,\dots,N\}$. In the function space, let $\mathcal{F}$ be the function class. The symbol $q_{\mathrm{data}}$ designates the target data distribution, while $p$ represents the conditional probability of LLM’s response (i.e., LLM policy).
@@ 2 Related Work
In the context of supervised fine-tuning (SFT) of LLMs, human-crafted data has proven to be a remarkably effective source that enhances the performance of LLMs on tasks such as code generation (Roziere et al., 2023; Yang et al., 2023) and mathematical reasoning (Yuan et al., 2023; Luo et al., 2023). While human data typically exhibits high quality, acquiring sufficient amount of such data poses a challenge in cost. In light of this consideration, the use of synthetic data has become increasingly popular and considered as a proxy for human data. This approach primarily leverages advanced LLMs such as the GPT series (Radford et al., 2019; Brown et al., 2020; OpenAI, 2023) as the guidance to generate high-quality data (Josifoski et al., 2023; Taori et al., 2023; Chiang et al., 2023; Li et al., 2023). Recent research has also highlighted the rephrasing capability of LLMs in prompting for better LLM response (Deng et al., 2023; Prasad et al., 2023) as well as augmenting synthetic data for more effective SFT (Yu et al., 2023; Liu et al., 2023). In contrast to prior studies that utilized more advanced models for synthetic data generation when pretraining or fine-tuning a target model, our approach directly generate synthetic data from the target model itself.
Curriculum Learning.
@@ 3.2 RL Fine-Tuning
$L_{\mathrm{RL}}(\bm{\theta})=\mathbb{E}_{\mathbf{x}\sim q(\cdot),\mathbf{y}\sim p_{\bm{\theta}}(\cdot|\mathbf{x})}[r(\mathbf{x},\mathbf{y})]-\lambda\mathbb{E}_{\mathbf{x}\sim q(\cdot)}\mathrm{KL}\big{(}p_{\bm{\theta}}(\cdot|\mathbf{x}))||p_{\mathrm{ref}}(\cdot|\mathbf{x})\big{)},$
@@
Meanwhile, the primary challenge in RL fine-tuning lies in finding a good reward function. Typically, this function requires training on a preference dataset. The compilation of such a dataset demands significant resources, often involving comprehensive evaluations either by human annotators, i.e., reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Bai et al., 2022a) or strong AI agents, i.e., reinforcement learning from AI feedback (RLAF) (Bai et al., 2022b).
@@ 4 Method
In this section, we introduce a new fine-tuning method for enhancing the performance of LLMs without relying on additional human or AI feedback. Consider a high-quality supervised fine-tuning (SFT) dataset $S_{\mathrm{SFT}}=\{(\mathbf{x},\mathbf{y})\}_{i=1}^{n}$, which are sampled from the marginal distribution $q(\mathbf{x})$ and $p_{\mathrm{data}}(\mathbf{y}|\mathbf{x})$. Given a supervised fine-tuned LLM $p_{\bm{\theta}_{0}}$, further application of the SFT approach in (3.1) with $S_{\mathrm{SFT}}$ will be ineffective and potentially lead to worse performance. In addition, without human and/or AI feedback, it becomes infeasible to acquire a preference dataset for RL fine-tuning (e.g., RLHF and RLAF). This hinders the application of RL fine-tuning techniques.
@@ 4.1 Self-Play Fine-Tuning (SPIN)
where $\ell(\cdot)$ is a loss function that is both monotonically decreasing and convex. For example, a linear loss function $\ell(t)=-t$ reduces (4.2) to the minimization version of (4.1). However, the use of a linear loss function results in an unbounded objective value, which, during continuous training, leads to a negative infinite value of $f(\mathbf{x},\mathbf{y}^{\prime})$ on the opponent player’s responses. Therefore, in our work, we choose the logistic loss function $\ell(t):=\log(1+\exp(-t))$ for its non-negativity, smoothness, and exponentially decaying tail as $t\rightarrow\infty$. Such a choice of loss function aids in preventing the excessive growth in the absolute value of $f$. It is worth noting that the objective function defined in (4.2) with a linear loss reduces to a similar IPM framework as in Wasserstein Generative Adversarial Networks (WGAN) (Arjovsky et al., 2017). However, our approach differs in both the choice of the function class $\mathcal{F}_{t}$ and the training procedure.
@@ 6.1 Experiment Setup
We employed the widely used Huggingface Open LLM Leaderboard (Beeching et al., 2023) as our evaluation benchmark, using the same Language Model Evaluation Harness library (Gao et al., 2023b). This leaderboard encompasses 6 different datasets, each focusing on a a specific capability of LLMs. Collectively, these datasets provide a thorough assessment framework, evaluating LLMs on commonsense reasoning (Arc (Clark et al., 2018), HellaSwag (Zellers et al., 2019), Winogrande (Sakaguchi et al., 2021)), multi-task language understanding (MMLU(Hendrycks et al., 2020)), human falsehood mimic (TruthfulQA (Lin et al., 2021)) and math problem solving (GSM8k (Cobbe et al., 2021)). In evaluation, the language models are prompted with few-shot in-context examples and the question. We follow the standard approach and report the average score across all datasets. In Table 1, we detail the evaluation setting adopted by both the leaderboard and our experiments. We leave further implementation details to Appendix A.
@@ 6.2 SPIN Effectively Improves Benchmark Performance
zephyr-7b-beta is a model derived from zephyr-7b-sft-full, trained with DPO on approximately $62$k preference data. This data, the UltraFeedback Binarized dataset [^2] https://huggingface.co/datasets/HuggingFaceH4/ultrafeedback_binarized, comprises both chosen and rejected completions evaluated by GPT-4. We note that, DPO requires either human input or advanced language model feedback to determine the preference, making data generation a rather expensive procedure. In contrast, our SPIN only requires the initial model itself. Moreover, unlike DPO which requires new data source, our method exclusively leverages the existing SFT dataset. In Figure 3, we show the performance comparison of SPIN at iterations 0 and 1 (employing $50$k SFT data) with DPO training, from the same SFT checkpoint. We can observe that, while DPO leverages more data from new sources, SPIN based on the existing SFT data can already achieve comparable average performance to DPO training at iteration 0. From iteration 1, SPIN even surpasses the performance of DPO on the leaderboard benchmark.
@@ 7 Conclusion and Discussion
Appendix A Experiment Details
A.1 Hyperparameters and Implementation Details
@@ A.2 Generation Examples
[…]
@@
How does the economic growth of Horsham compare to nearby towns?
@@
Appendix B Proof of Theorems in Section 5
B.1 Proof of Theorem 5.2
@@ Appendix B Proof of Theorems in Section 5
Lemma B.1 shows that the global minimum of $a\ell(t)+b\ell(-t)$ is achieved when $t=\log(a/b)$. Based on Lemma B.1, we can further prove that (4.2) with the logistic loss function has a closed-form solution if we ignore the constraint set $\mathcal{F}_{t}$.
Lemma B.2.
@@
where the inequality is due to $a\ell(t)+b\ell(-t)\geq a\log(1+b/a)+b\log(1+a/b)$ in Lemma B.1 with $a=q(\mathbf{x})p_{\mathrm{data}}(\mathbf{y}|\mathbf{x})p_{\bm{\theta}_{t}}(\mathbf{y}^{\prime}|\mathbf{x}),b=q(\mathbf{x})p_{\mathrm{data}}(\mathbf{y}^{\prime}|\mathbf{x})p_{\bm{\theta}_{t}}(\mathbf{y}|\mathbf{x})$, $t=f(\mathbf{x},\mathbf{y})-f(\mathbf{x},\mathbf{y}^{\prime})$. The equality (i) holds if and only if the following equation holds almost surely for any $\mathbf{x},\mathbf{y},\mathbf{y}^{\prime}$,
$f(\mathbf{x},\mathbf{y})-f(\mathbf{x},\mathbf{y}^{\prime})=\log\bigg{(}\frac{p_{\mathrm{data}}(\mathbf{y}|\mathbf{x})p_{\bm{\theta}_{t}}(\mathbf{y}^{\prime}|\mathbf{x})}{p_{\mathrm{data}}(\mathbf{y}^{\prime}|\mathbf{x})p_{\bm{\theta}_{t}}(\mathbf{y}|\mathbf{x})}\bigg{)}.$ (B.1)
(B.1) is equivalent to
@@
Lemma B.2 provides a closed-form solution to (4.2) if we ignore the constraint set $\mathcal{F}_{t}$. If this closed-form solution belongs to $\mathcal{F}_{t}$, then it should also be the solution to (4.2). This observation is the key to the proof of Theorem 5.4.
@@
$p_{\bm{\theta}}(\mathbf{y}|\mathbf{x})=\widehat{Z}(\mathbf{x})\cdot p_{\bm{\theta}_{t}}(\mathbf{y}|\mathbf{x})\big{(}p_{\mathrm{data}}(\mathbf{y}|\mathbf{x})/p_{\bm{\theta}_{t}}(\mathbf{y}|\mathbf{x})\big{)}^{1/\lambda}.$ (B.2)
Applying logarithm function on both side of (B.2) yields
@@
By Lemma B.2, $f^{*}(\mathbf{x},\mathbf{y})=\lambda\log(\widehat{Z}(\mathbf{x}))+\log\big{(}\frac{p_{\mathrm{data}}(\mathbf{y}|\mathbf{x})}{p_{\bm{\theta}_{t}}(\mathbf{y}|\mathbf{x})}\big{)}$ is the global minimum of the following minimization problem,
$\mathop{\mathrm{argmin}}_{f}\mathbb{E}_{\mathbf{y}\sim p_{\mathrm{data}}(\cdot|\mathbf{x}),y^{\prime}\sim p_{\bm{\theta}_{t}}(\cdot|\mathbf{x})}\big{[}\ell\big{(}f(\mathbf{x},\mathbf{y})-f(\mathbf{x},\mathbf{y}^{\prime})\big{)}\big{]}.$ (B.3)
@@
$=\min_{\bm{\theta}\in\bm{\Theta}}L_{\texttt{SPIN}}(\bm{\theta},\bm{\theta}_{t}).$ (B.4)
Since $\bm{\theta}_{t+1}$ is the global minimum of $L_{\texttt{SPIN}}(\bm{\theta},\bm{\theta}_{t})$. Then by (B.4), $\lambda\log\Big{(}\frac{p_{\bm{\theta}_{t+1}}(\mathbf{y}|\mathbf{x})}{p_{\bm{\theta}_{t}}(\mathbf{y}|\mathbf{x})}\Big{)}$ should be the global minimum of problem (B.3). By Lemma B.2, there exists $Z(\mathbf{x})$ such that

### 12 Feb 2024 v2

@@ Abstract
Harnessing the power of human-annotated data through Supervised Fine-Tuning (SFT) is pivotal for advancing Large Language Models (LLMs). In this paper, we delve into the prospect of growing a strong LLM out of a weak one without the need for acquiring additional human-annotated data. We propose a new fine-tuning method called Self-Play fIne-tuNing (SPIN), which starts from a supervised fine-tuned model. At the heart of SPIN lies a self-play mechanism, where the LLM refines its capability by playing against instances of itself. More specifically, the LLM generates its own training data from its previous iterations, refining its policy by discerning these self-generated responses from those obtained from human-annotated data. Our method progressively elevates the LLM from a nascent model to a formidable one, unlocking the full potential of human-annotated demonstration data for SFT. Theoretically, we prove that the global optimum to the training objective function of our method is achieved only when the LLM policy aligns with the target data distribution. Empirically, we evaluate our method on several benchmark datasets including the HuggingFace Open LLM Leaderboard, MT-Bench, and datasets from Big-Bench. Our results show that SPIN can significantly improve the LLM’s performance across a variety of benchmarks and even outperform models trained through direct preference optimization (DPO) supplemented with extra GPT-4 preference data. This sheds light on the promise of self-play, enabling the achievement of human-level performance in LLMs without the need for expert opponents. Codes are available at https://github.com/uclaml/SPIN.
@@ 1 Introduction
Notation.
We use lowercase letters and lowercase boldface letters to denote scalars and vectors, respectively. We use $[N]$ to denote the index set $\{1,\dots,N\}$. In the function space, let $\mathcal{F}$ be the function class. The symbol $q_{\mathrm{data}}$ designates the target data distribution, while $p$ represents the conditional probability of LLM’s response (i.e., LLM policy).
@@ 2 Related Work
In the context of supervised fine-tuning (SFT) of LLMs, human-crafted data has proven to be a remarkably effective source that enhances the performance of LLMs on tasks such as code generation (Roziere et al., 2023; Yang et al., 2023) and mathematical reasoning (Yuan et al., 2023; Luo et al., 2023). While human data typically exhibits high quality, acquiring sufficient amount of such data poses a challenge in cost. In light of this consideration, the use of synthetic data has become increasingly popular and considered as a proxy for human data. This approach primarily leverages advanced LLMs such as the GPT series (Radford et al., 2019; Brown et al., 2020; OpenAI, 2023) as the guidance to generate high-quality data (Josifoski et al., 2023; Taori et al., 2023; Chiang et al., 2023; Li et al., 2023). Recent research has also highlighted the rephrasing capability of LLMs in prompting for better LLM response (Deng et al., 2023; Prasad et al., 2023) as well as augmenting synthetic data for more effective SFT (Yu et al., 2023; Liu et al., 2023). In contrast to prior studies that utilized more advanced models for synthetic data generation when pre-training or fine-tuning a target model, our approach directly generates synthetic data from the target model itself.
@@ 3.2 RL Fine-Tuning
$L_{\mathrm{RL}}(\bm{\theta})=\mathbb{E}_{\mathbf{x}\sim q(\cdot),\mathbf{y}\sim p_{\bm{\theta}}(\cdot|\mathbf{x})}[r(\mathbf{x},\mathbf{y})]-\lambda\mathbb{E}_{\mathbf{x}\sim q(\cdot)}\mathrm{KL}\big{(}p_{\bm{\theta}}(\cdot|\mathbf{x})||p_{\mathrm{ref}}(\cdot|\mathbf{x})\big{)},$
@@
Meanwhile, the primary challenge in RL fine-tuning lies in finding a good reward function. Typically, this function requires training on a preference dataset. The compilation of such a dataset demands significant resources, often involving comprehensive evaluations either by human annotators, i.e., reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Bai et al., 2022a) or strong AI agents, i.e., reinforcement learning from AI feedback (RLAIF) (Bai et al., 2022b).
@@ 4 Method
In this section, we introduce a new fine-tuning method for enhancing the performance of LLMs without relying on additional human or AI feedback. Consider a high-quality supervised fine-tuning (SFT) dataset $S_{\mathrm{SFT}}=\{(\mathbf{x},\mathbf{y})\}_{i=1}^{n}$, which are sampled from the marginal distribution $q(\mathbf{x})$ and $p_{\mathrm{data}}(\mathbf{y}|\mathbf{x})$. Given a supervised fine-tuned LLM $p_{\bm{\theta}_{0}}$, further application of the SFT approach in (3.1) with $S_{\mathrm{SFT}}$ will be ineffective and potentially lead to worse performance. In addition, without human and/or AI feedback, it becomes infeasible to acquire a preference dataset for RL fine-tuning (e.g., RLHF and RLAIF). This hinders the application of RL fine-tuning techniques.
@@ 4.1 Self-Play Fine-Tuning (SPIN)
where $\ell(\cdot)$ is a loss function that is both monotonically decreasing and convex. For example, a linear loss function $\ell(t)=-t$ reduces (4.2) to the minimization version of (4.1). However, the use of a linear loss function results in an unbounded objective value, which, during continuous training, leads to a negative infinite value of $f(\mathbf{x},\mathbf{y}^{\prime})$ on the opponent player’s responses. Therefore, in our work, we choose the logistic loss function $\ell(t):=\log(1+\exp(-t))$ for its non-negativity, smoothness, and exponentially decaying tail as $t\rightarrow\infty$. Such a choice of loss function aids in preventing the excessive growth in the absolute value of $f$.
@@ 6.1 Experiment Setup
We employed the widely used Huggingface Open LLM Leaderboard (Beeching et al., 2023) as our evaluation benchmark, using the same Language Model Evaluation Harness library (Gao et al., 2023b). This leaderboard encompasses 6 different datasets, each focusing on a a specific capability of LLMs. Collectively, these datasets provide a thorough assessment framework, evaluating LLMs on commonsense reasoning (Arc (Clark et al., 2018), HellaSwag (Zellers et al., 2019), Winogrande (Sakaguchi et al., 2021)), multi-task language understanding (MMLU(Hendrycks et al., 2020)), human falsehood mimic (TruthfulQA (Lin et al., 2021)) and math problem solving (GSM8k (Cobbe et al., 2021)). In evaluation, the language models are prompted with few-shot in-context examples and the question. We follow the standard approach and report the average score across all datasets. In Table 1, we detail the evaluation setting adopted by both the leaderboard and our experiments. We leave further implementation details to Appendix B.
@@ 6.2 SPIN Effectively Improves Benchmark Performance
zephyr-7b-beta is a model derived from zephyr-7b-sft-full, trained with DPO on approximately $62$k preference data. This data, the UltraFeedback Binarized dataset(Cui et al., 2023) [^2] https://huggingface.co/datasets/HuggingFaceH4/ultrafeedback_binarized, comprises both chosen and rejected completions evaluated by GPT-4. We note that, DPO requires either human input or advanced language model feedback to determine the preference, making data generation a rather expensive procedure. In contrast, our SPIN only requires the initial model itself. Moreover, unlike DPO which requires new data source, our method exclusively leverages the existing SFT dataset. In Figure 3, we show the performance comparison of SPIN at iterations 0 and 1 (employing $50$k SFT data) with DPO training, from the same SFT checkpoint. We can observe that, while DPO leverages more data from new sources, SPIN based on the existing SFT data can already achieve comparable average performance to DPO training at iteration 0. From iteration 1, SPIN even surpasses the performance of DPO on the leaderboard benchmark.
@@ 7 Conclusion and Discussion
Appendix A Further Related Work
Curriculum Learning.
Generative Adversarial Learning.
Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) represent a distinct class of generative models, characterized by their unique adversarial process. To enhance training stability and data quality, Mao et al. (2017) introduced the Least Squares GAN, employing a least squares loss function for the discriminator. A significant advancement in GANs involves the use of Integral Probability Metrics (IPM) (Müller, 1997), particularly highlighted in the development of Wasserstein GAN by Arjovsky et al. (2017). This model employs IPM in its loss design, enhancing training stability. Since then, IPMs have become popular in the design of GANs (Mroueh and Sercu, 2017; Gulrajani et al., 2017), particularly in constraining the discriminator to a specific function class, thereby preventing it from overpowering the generator. Furthermore, Jolicoeur-Martineau (2018) generalized IPM-based GANs by introducing relativistic discriminator and proposed Relativistic GAN. It is worth noting that the objective function defined in our (4.2) is similar to Relativistic GAN (Jolicoeur-Martineau, 2018) and reduces to an IPM framework in Wasserstein GAN (Arjovsky et al., 2017) with a linear loss. However, our approach differs in both the choice of the function class and the training procedure. Inspired by GAN, Cheng et al. (2023) proposed an adversarial learning framework named Adversarial Preference Optimization (APO) that trains the LLM and a reward model in an adversarial game. Our method is also related to Generative Adversarial Imitation Learning (GAIL) (Ho and Ermon, 2016), which trains separate discriminator and policy networks in each iteration for imitation learning. In contrast to the above methods, SPIN relies on self-play where both the main player and the opponent player are the same LLM from two consecutive iterations.
Appendix B Experiment Details
B.1 Hyperparameters and Implementation Details
@@ B.2 Generation Examples
How does the economic growth of Horsham compare to nearby towns?
@@
[…]
@@
Appendix C Proof of Theorems in Section 5
C.1 Proof of Theorem 5.2
@@ Appendix C Proof of Theorems in Section 5
Lemma C.1 shows that the global minimum of $a\ell(t)+b\ell(-t)$ is achieved when $t=\log(a/b)$. Based on Lemma C.1, we can further prove that (4.2) with the logistic loss function has a closed-form solution if we ignore the constraint set $\mathcal{F}_{t}$.
Lemma C.2.
@@
where the inequality is due to $a\ell(t)+b\ell(-t)\geq a\log(1+b/a)+b\log(1+a/b)$ in Lemma C.1 with $a=q(\mathbf{x})p_{\mathrm{data}}(\mathbf{y}|\mathbf{x})p_{\bm{\theta}_{t}}(\mathbf{y}^{\prime}|\mathbf{x}),b=q(\mathbf{x})p_{\mathrm{data}}(\mathbf{y}^{\prime}|\mathbf{x})p_{\bm{\theta}_{t}}(\mathbf{y}|\mathbf{x})$, $t=f(\mathbf{x},\mathbf{y})-f(\mathbf{x},\mathbf{y}^{\prime})$. The equality (i) holds if and only if the following equation holds almost surely for any $\mathbf{x},\mathbf{y},\mathbf{y}^{\prime}$,
$f(\mathbf{x},\mathbf{y})-f(\mathbf{x},\mathbf{y}^{\prime})=\log\bigg{(}\frac{p_{\mathrm{data}}(\mathbf{y}|\mathbf{x})p_{\bm{\theta}_{t}}(\mathbf{y}^{\prime}|\mathbf{x})}{p_{\mathrm{data}}(\mathbf{y}^{\prime}|\mathbf{x})p_{\bm{\theta}_{t}}(\mathbf{y}|\mathbf{x})}\bigg{)}.$ (C.1)
Equation (C.1) is equivalent to
@@
Lemma C.2 provides a closed-form solution to (4.2) if we ignore the constraint set $\mathcal{F}_{t}$. If this closed-form solution belongs to $\mathcal{F}_{t}$, then it should also be the solution to (4.2). This observation is the key to the proof of Theorem 5.4.
@@
$p_{\bm{\theta}}(\mathbf{y}|\mathbf{x})=\widehat{Z}(\mathbf{x})\cdot p_{\bm{\theta}_{t}}(\mathbf{y}|\mathbf{x})\big{(}p_{\mathrm{data}}(\mathbf{y}|\mathbf{x})/p_{\bm{\theta}_{t}}(\mathbf{y}|\mathbf{x})\big{)}^{1/\lambda}.$ (C.2)
Applying logarithm function on both side of (C.2) yields
@@
By Lemma C.2, $f^{*}(\mathbf{x},\mathbf{y})=\lambda\log(\widehat{Z}(\mathbf{x}))+\log\big{(}\frac{p_{\mathrm{data}}(\mathbf{y}|\mathbf{x})}{p_{\bm{\theta}_{t}}(\mathbf{y}|\mathbf{x})}\big{)}$ is the global minimum of the following minimization problem,
$\mathop{\mathrm{argmin}}_{f}\mathbb{E}_{\mathbf{y}\sim p_{\mathrm{data}}(\cdot|\mathbf{x}),y^{\prime}\sim p_{\bm{\theta}_{t}}(\cdot|\mathbf{x})}\big{[}\ell\big{(}f(\mathbf{x},\mathbf{y})-f(\mathbf{x},\mathbf{y}^{\prime})\big{)}\big{]}.$ (C.3)
@@
$=\min_{\bm{\theta}\in\bm{\Theta}}L_{\texttt{SPIN}}(\bm{\theta},\bm{\theta}_{t}).$ (C.4)
Since $\bm{\theta}_{t+1}$ is the global minimum of $L_{\texttt{SPIN}}(\bm{\theta},\bm{\theta}_{t})$. Then by (C.4), $\lambda\log\Big{(}\frac{p_{\bm{\theta}_{t+1}}(\mathbf{y}|\mathbf{x})}{p_{\bm{\theta}_{t}}(\mathbf{y}|\mathbf{x})}\Big{)}$ should be the global minimum of problem (C.3). By Lemma C.2, there exists $Z(\mathbf{x})$ such that